# Drell–Yan at the Electron-Ion Collider

### v1

@@ III.2 Spin Asymmetries
The projections for the measurement of the target-spin asymmetries as a function of pair $p_{T}$ for several bins of $M_{\ell\ell}$ are shown in Fig. 11. The proton beam polarization is taken to be $70\pm 1\%$. The total uncertainty in Fig. 11 includes the polarization uncertainty, the uncertainty on the asymmetry of the background, and the statistical uncertainty. The results remain the same whether the target is polarized longitudinally or transversely, assuming the same level of uncertainty can be reached on the asymmetries of the backgrounds. Fig. 11 shows that the uncertainty on the asymmetry at low $M_{\ell\ell}$ is large and dominated by the uncertainty from the background while the large $M_{\ell\ell}$ data is reasonably precise since the amount of background is expected to be lower. It should be noted that the $y$-axis scale in Fig. 11 is very large, the uncertainty on the most precise asymmetry points, i.e. those at large $M_{\ell\ell}$ and low $p_{T,\ell\ell}$, is around 5%.
@@ IV.1 Charge-odd lepton asymmetry from Bethe–Heitler–Drell–Yan interference
Instead of serving merely as a background to the Drell–Yan signal, the Bethe–Heitler process can in principle be turned into a tool to access novel combinations of photon and proton parton distributions, via its interference with partonic amplitudes. An analogous situation occurs in DVCS and TCS, where the interference between Bethe–Heitler and the partonic Compton amplitude provides sensitivity to the real and imaginary parts of the Compton form factors. We focus first on the interference between Bethe–Heitler and the standard Drell–Yan mechanism in unpolarized $\gamma p\to\ell^{+}\ell^{-}X$, following the setup of Ref. Busenitz:1981hr.
The interference between the Bethe–Heitler and Drell–Yan processes is largest in the two–jet topology. For Bethe–Heitler, this corresponds to the process in the top left panel of Fig. 12, where the exchanged photon between the lepton and the proton carries sizable momentum and produces a target jet and a current jet in the final state. In this configuration the Bethe–Heitler process becomes experimentally indistinguishable, at the level of hadronic activity, from the direct Drell–Yan contribution shown in the top right panel of Fig. 12, in which the incoming photon couples directly to a quark without leaving a spectator remnant in the photon direction. It should be reiterated that this contribution from direct photoproduction is not included in the Sherpa simulation discussed in the rest of this paper. However, the ratio of the cross sections for direct and resolved photoproduction of jets at HERA was observed to be of order unity, indicating that the direct yields should be within an order of magnitude of the resolved yields obtained from Sherpa ZEUS:2001zoq; H1:2002apm; ZEUS:1995xfc; H1:2006rre. Resolved photoproduction will also interfere with Bethe–Heitler, but to a lesser degree, since the kinematics of the jets and leptons in resolved photoproduction at $x_{\gamma}\ll 1$ are significantly different from those at $x_{\gamma}\approx 1$.
@@
To provide sensitivity to the Bethe–Heitler–Drell–Yan interference term, one can define an interference–driven angular asymmetry in the dilepton rest frame that is identically zero for the pure Bethe–Heitler and pure Drell–Yan contributions, and nonzero only for their interference. We work in the $\ell^{+}\ell^{-}$ center-of-mass frame and specify the vector of the positively charged lepton via polar and azimuthal angles $(\theta,\phi)$ with respect to the incoming-photon direction, with the $xz$ plane chosen to contain the proton momentum. Under interchange of the lepton charge one has $(\theta,\phi)\to(\pi-\theta,\phi+\pi)$; the Bethe–Heitler and Drell–Yan squared amplitudes are even under this transformation, whereas the interference term is odd. To define an asymmetry sensitive to the interference, the lepton phase space can be divided into two azimuthal hemispheres, denoted here as left and right, defined by $\phi\in[\pi/2,3\pi/2]$ and $\phi\in[-\pi/2,\pi/2]$, respectively, integrating over the available range in $\cos\theta$. An asymmetry can then be written simply as
@@
At the EIC the situation for resolving these interferences is favorable in several respects. Resolved photoproduction Drell–Yan cross sections are at the few-pb level, implying on the order of $10^{5}$ direct photoproduction dilepton events in $100~\mathrm{fb}^{-1}$. In the resolved photoproduction analysis above we studied regions in $x_{\gamma}$, where the Bethe–Heitler contribution is suppressed and the resolved-photon Drell–Yan signal dominates. The interference term also vanishes in that case when integrating over $\theta$ and $\phi$. However, one can deliberately select the complementary kinematic regime where Bethe–Heitler becomes large and the interfering topology is realized, i.e. $x_{\gamma}\to 1$ with two hard jets. In this region, the high luminosity and large acceptance at the EIC make it feasible to measure charge-odd observables such as $A_{\rm INT}$. The limitation on such an asymmetry measurement will likely be the understanding of the background asymmetry that must be subtracted, similar to the target-spin asymmetry case discussed previously.
@@ IV.2 Access to transversity
In addition to the unpolarized charge asymmetry $A_{\rm INT}$, transverse target-spin asymmetries have been predicted to be sensitive to the nucleon’s transversity distribution and to the chiral–odd distribution amplitude of the real photon Pire:2009ap. Extractions of the quark transversity distribution in semi-inclusive DIS require coupling the transversity PDF $h_{1}^{q}(x)$ to a chiral-odd fragmentation function. In the single-hadron channel this role is played by the Collins fragmentation function $H_{1}^{\perp,q}(z,p_{T}^{2})$, taken from global fits constrained by SIDIS and $e^{+}e^{-}$ data from HERMES, COMPASS, and Belle Airapetian:2010ds; Adolph:2012sn; Seidl:2008xc; Anselmino:2007fs; Anselmino:2013vqa. An alternative method uses dihadron SIDIS and the chiral-odd interference fragmentation function $H_{1}^{\sphericalangle,q}(z,M_{h}^{2})$. At the EIC, where the transversity PDF is a key physics deliverable, it is desirable to have a way to cross check whether the measured SIDIS asymmetries are a result of the fragmentation functions or the PDFs. Ref. Pire:2009ap proposes the dilepton photoproduction process as a novel avenue for doing precisely this.
@@
The arguments of $\phi_{\gamma}(z)$ and $h^{q}_{1}$ are suppressed for clarity, the reader should refer to Ref. Pire:2009ap for the details. The magnetic susceptibility and photon distribution amplitude can be extracted from e.g. QCD sum rules or lattice simulations and used as inputs to relate the asymmetry directly to $h^{q}_{1}$. Alternatively, one can take the transversity from SIDIS as an input and attempt to extract the photon distribution amplitude or QCD magnetic susceptibility. This cross section difference can also be turned into an asymmetry for easier experimental measurement.
The highly polarized proton and electron beams at the EIC compel the pursuit of this transversity program in detail. The particle identification capability of the EIC also opens the door for identifying the flavor of the struck quark, providing some access to the flavor-dependent transversity. A full experimental study of the precision on the transversity PDF enabled by this process the EIC is beyond the scope of this work, and will rely on an actual numerical estimate for the cross section difference $\Delta\sigma_{T}$ in the EIC phase space. Nevertheless, the EIC provides all the tools necessary to make such measurements of transversity, provided the numerical size of the cross section difference is not vanishingly small. While the extraction of transversity will rely on additional non-perturbative inputs, i.e. the photon distribution amplitude and QCD magnetic susceptibility, the lack of dependence on the fragmentation functions breaks the degeneracy that exists in SIDIS, thereby permitting more rigorous extraction of the transversity PDFs from the EIC data if this contribution can be measured.
@@ V Pion-photon Drell–Yan
Another possibility offered by the EIC is to study Drell–Yan on the pion in leading-neutron photoproduction, i.e. the Sullivan process Sullivan:1971kd [^9] A similar measurement was recently discussed in the context of ultra-peripheral collisions at the LHC Goncalves:2025aay, although that study focuses on accessing the pion gluon distribution via heavy quark photoproduction.. In events where a neutron carries a large fraction of the incoming proton beam energy $x_{L}=E_{n}/E_{p}\gtrsim 0.8$ and has a small transverse momentum, the HERA data support the so-called proton vertex factorization, whereby the cross section factorizes into a pion flux factor and the hard scattering on a parton inside the pion. Therefore, in the kinematic regime where proton vertex factorization applies, the photoproduction of dileptons and a leading neutron can be interpreted as Drell–Yan between the photon and a $\pi^{+}$ Sullivan:1971kd; Kopeliovich:1996iw. The leading-order diagram of this process is shown in Fig. 13.
Leading-neutron DIS data from HERA, i.e. $ep\rightarrow e^{\prime}n+X$, provides the core dataset for extractions of the parton distributions of the pion ZEUS:1996uej; ZEUS:2002gig; H1:2010hym; Kopeliovich:2012fd; ZEUS:2007knd; McKenney:2015xis; Barry:2018ort; Novikov:2020snp; Cao:2021aci; Barry:2025wjx. At HERA high-energy leading neutrons were observed in 1-2% of inclusive DIS events ZEUS:2002gig and 5% of all dijet photoproduction events ZEUS:2000qdw; a similar percent-level fraction can be expected at the EIC. The one-pion-exchange model describes the ZEUS data well when the leading neutron carries 60% or more of the incoming proton beam energy ZEUS:2002gig. Therefore, for pions carrying less than 40% of the incoming proton beam energy, proton vertex factorization seems to hold. JAM imposes a stricter requirement of $x_{L}>0.8$, corresponding to pions with less than 20% of the incoming proton beam energy. Applying this requirement, the accessible pion energies for the 10x275 GeV configuration are only 55 GeV and below. It should be remembered that only the region of $1\lesssim E_{\gamma}\lesssim 6~\mathrm{GeV}$ resides within the Low-$Q^{2}$ tagger acceptance, so the range of center-of-mass energies is only roughly $7\lesssim\sqrt{s_{\gamma\pi}}\lesssim 36\mathrm{~GeV}$. This range of energies is similar to existing and proposed fixed-target Drell–Yan experiments at Fermilab, CERN, and JPARC.
@@
Figure 13: Diagram of leading order Drell–Yan in $\gamma\pi^{+}$. Accessible at the EIC in the case of a leading neutron with a rapidity gap between the neutron and the pion remnant state $X_{\pi}$.
@@
The primary benefits of the $\gamma\pi$ Drell–Yan measurement at the EIC are that it can serve as a test for various assumptions and that novel regions of phase space in $Q^{2}$ and $x_{\pi}$ may be accessible. It is unlikely that the measurement precision at the EIC will compete with the precision of the leading-neutron DIS or pion Drell–Yan data. However, the universality of proton vertex factorization can be tested for the first time in a process other than leading-neutron DIS, helping to validate the assumption that pion structure is accessible in leading-neutron processes.
@@ VII Acknowledgments
I am grateful to the Sherpa team for implementing resolved photoproduction in their event generator, and P. Meinzinger in particular for help understanding and running Sherpa. Without their efforts this work would not have been possible.
@@ Appendix A Detector Resolutions and Acceptances
The strawman detector used for the reconstruction studies follows closely the design of ePIC. When handling charged particles, whichever detector resolution is best is the assumed resolution, i.e. for pions in the barrel the tracking resolution is used over the hadronic calorimetry resolution. $K^{0}_{s}$ and $\Lambda$ are assumed to have the momentum resolution of the tracking. Neutrons and $K^{0}_{L}$ are produced in roughly equal numbers, so long-lived neutral hadrons are assigned the average of the two masses, i.e. 0.72 GeV. Particle identification of charged hadrons is assumed to be correct within the central detector acceptance.
@@
Table 4: Low-$Q^{2}$ tagger smearing and acceptance applied to the scattered electron in the tagged sample.

### v2

@@ III.2 Spin Asymmetries
The projections for the measurement of the target-spin asymmetries as a function of pair $p_{T}$ for several bins of $M_{\ell\ell}$ are shown in Fig. 11. The proton beam polarization is taken to be $70\pm 1\%$. The total uncertainty in Fig. 11 includes the polarization uncertainty, the uncertainty on the asymmetry of the background, and the statistical uncertainty. The results remain the same whether the target is polarized longitudinally or transversely, assuming the same level of uncertainty can be reached on the asymmetries of the backgrounds. Fig. 11 shows that the uncertainty on the asymmetry at low $M_{\ell\ell}$ is large and dominated by the uncertainty from the background while the large $M_{\ell\ell}$ data is reasonably precise since the amount of background is expected to be lower. It should be noted that the $y$-axis scale in Fig. 11 is very large: the uncertainty on the most precise asymmetry points, i.e., those at large $M_{\ell\ell}$ and low $p_{T,\ell\ell}$, is around 5%.
@@ IV.1 Charge-odd lepton asymmetry from Bethe–Heitler–Drell–Yan interference
Instead of serving merely as a background to the Drell–Yan signal, the Bethe–Heitler process can in principle be turned into a tool to access novel combinations of photon and proton parton distributions via its interference with partonic amplitudes. An analogous situation occurs in DVCS and TCS, where the interference between Bethe–Heitler and the partonic Compton amplitude provides sensitivity to the real and imaginary parts of the Compton form factors. We focus first on the interference between Bethe–Heitler and the standard Drell–Yan mechanism in unpolarized $\gamma p\to\ell^{+}\ell^{-}X$, following the setup of Ref. Busenitz:1981hr.
The interference between the Bethe–Heitler and Drell–Yan processes is largest in the two–jet topology. For Bethe–Heitler, this corresponds to the process in the top left panel of Fig. 12, where the exchanged photon between the lepton and the proton carries sizable momentum and produces a target jet and a current jet in the final state. In this configuration, the Bethe–Heitler process becomes experimentally indistinguishable, at the level of hadronic activity, from the direct Drell–Yan contribution shown in the top right panel of Fig. 12, in which the incoming photon couples directly to a quark without leaving a spectator remnant in the photon direction. It should be reiterated that this contribution from direct photoproduction is not included in the Sherpa simulation discussed in the rest of this paper. However, the ratio of the cross sections for direct and resolved photoproduction of jets at HERA was observed to be of order unity, indicating that the direct yields should be within an order of magnitude of the resolved yields obtained from Sherpa ZEUS:2001zoq; H1:2002apm; ZEUS:1995xfc; H1:2006rre. Resolved photoproduction will also interfere with Bethe–Heitler, but to a lesser degree, since the kinematics of the jets and leptons in resolved photoproduction at $x_{\gamma}\ll 1$ are significantly different from those at $x_{\gamma}\approx 1$.
@@
To provide sensitivity to the Bethe–Heitler–Drell–Yan interference term, one can define an interference–driven angular asymmetry in the dilepton rest frame that is identically zero for the pure Bethe–Heitler and pure Drell–Yan contributions and nonzero only for their interference. We work in the $\ell^{+}\ell^{-}$ center-of-mass frame and specify the vector of the positively charged lepton via polar and azimuthal angles $(\theta,\phi)$ with respect to the incoming-photon direction, with the $xz$ plane chosen to contain the proton momentum. Under interchange of the lepton charge one has $(\theta,\phi)\to(\pi-\theta,\phi+\pi)$; the Bethe–Heitler and Drell–Yan squared amplitudes are even under this transformation, whereas the interference term is odd. To define an asymmetry sensitive to the interference, the lepton phase space can be divided into two azimuthal hemispheres, denoted here as left and right, defined by $\phi\in[\pi/2,3\pi/2]$ and $\phi\in[-\pi/2,\pi/2]$, respectively, integrating over the available range in $\cos\theta$. An asymmetry can then be written simply as
@@
At the EIC, the situation for resolving these interferences is favorable in several respects. Resolved photoproduction Drell–Yan cross sections are at the few-pb level, implying on the order of $10^{5}$ direct photoproduction dilepton events in $100~\mathrm{fb}^{-1}$. In the resolved photoproduction analysis above, we studied regions in $x_{\gamma}$, where the Bethe–Heitler contribution is suppressed and the resolved-photon Drell–Yan signal dominates. The interference term also vanishes in that case when integrating over $\theta$ and $\phi$. However, one can deliberately select the complementary kinematic regime where Bethe–Heitler becomes large and the interfering topology is realized, i.e., $x_{\gamma}\to 1$ with two hard jets. In this region, the high luminosity and large acceptance at the EIC make it feasible to measure charge-odd observables such as $A_{\rm INT}$. The limitation on such an asymmetry measurement will likely be the understanding of the background asymmetry that must be subtracted, similar to the target-spin asymmetry case discussed previously.
@@ IV.2 Access to transversity
In addition to the unpolarized charge asymmetry $A_{\rm INT}$, transverse target-spin asymmetries have been predicted to be sensitive to the nucleon’s transversity distribution and to the chiral–odd distribution amplitude of the real photon Pire:2009ap. Extractions of the quark transversity distribution in semi-inclusive DIS require coupling the transversity PDF $h_{1}^{q}(x)$ to a chiral-odd fragmentation function. In the single-hadron channel this role is played by the Collins fragmentation function $H_{1}^{\perp,q}(z,p_{T}^{2})$, taken from global fits constrained by SIDIS and $e^{+}e^{-}$ data from HERMES, COMPASS, and Belle Airapetian:2010ds; Adolph:2012sn; Seidl:2008xc; Anselmino:2007fs; Anselmino:2013vqa; DAlesio:2020vtw. An alternative method uses dihadron SIDIS and the chiral-odd interference fragmentation function $H_{1}^{\sphericalangle,q}(z,M_{h}^{2})$. At the EIC, where the transversity PDF is a key physics deliverable, it is desirable to have a way to cross check whether the measured SIDIS asymmetries are a result of the fragmentation functions or the PDFs. Ref. Pire:2009ap proposes the dilepton photoproduction process as a novel avenue for doing precisely this.
@@
The arguments of $\phi_{\gamma}(z)$ and $h^{q}_{1}$ are suppressed for clarity; the reader should refer to Ref. Pire:2009ap for the details. The magnetic susceptibility and photon distribution amplitude can be extracted from, e.g., QCD sum rules or lattice simulations and used as inputs to relate the asymmetry directly to $h^{q}_{1}$. Alternatively, one can take the transversity from SIDIS as an input and attempt to extract the photon distribution amplitude or QCD magnetic susceptibility. This cross section difference can also be turned into an asymmetry for easier experimental measurement.
The highly polarized proton and electron beams at the EIC compel the pursuit of this transversity program in detail. The particle identification capability of the EIC also opens the door for identifying the flavor of the struck quark, providing some access to the flavor-dependent transversity. A full experimental study of the precision on the transversity PDF enabled by this process at the EIC is beyond the scope of this work and will rely on an actual numerical estimate for the cross section difference $\Delta\sigma_{T}$ in the EIC phase space. Nevertheless, the EIC provides all the tools necessary to make such measurements of transversity, provided the numerical size of the cross section difference is not vanishingly small. While the extraction of transversity will rely on additional non-perturbative inputs, i.e., the photon distribution amplitude and QCD magnetic susceptibility, the lack of dependence on the fragmentation functions breaks the degeneracy that exists in SIDIS, thereby permitting more rigorous extraction of the transversity PDFs from the EIC data if this contribution can be measured.
@@ V Pion-photon Drell–Yan
Another possibility offered by the EIC is to study Drell–Yan on the pion in leading-neutron photoproduction, i.e., the Sullivan process Sullivan:1971kd [^9] A similar measurement was recently discussed in the context of ultra-peripheral collisions at the LHC Goncalves:2025aay, although that study focuses on accessing the pion gluon distribution via heavy quark photoproduction.. In events where a neutron carries a large fraction of the incoming proton beam energy $x_{L}=E_{n}/E_{p}\gtrsim 0.8$ and has a small transverse momentum, the HERA data support the so-called proton vertex factorization, whereby the cross section factorizes into a pion flux factor and the hard scattering on a parton inside the pion. Therefore, in the kinematic regime where proton vertex factorization applies, the photoproduction of dileptons and a leading neutron can be interpreted as Drell–Yan between the photon and a $\pi^{+}$ Sullivan:1971kd; Kopeliovich:1996iw. The leading-order diagram of this process is shown in Fig. 13.
Leading-neutron DIS data from HERA, i.e., $ep\rightarrow e^{\prime}n+X$, provides the core dataset for extractions of the parton distributions of the pion ZEUS:1996uej; ZEUS:2002gig; H1:2010hym; Kopeliovich:2012fd; ZEUS:2007knd; McKenney:2015xis; Barry:2018ort; Novikov:2020snp; Cao:2021aci; Barry:2025wjx. At HERA high-energy leading neutrons were observed in 1-2% of inclusive DIS events ZEUS:2002gig and 5% of all dijet photoproduction events ZEUS:2000qdw; a similar percent-level fraction can be expected at the EIC. The one-pion-exchange model describes the ZEUS data well when the leading neutron carries 60% or more of the incoming proton beam energy ZEUS:2002gig. Therefore, for pions carrying less than 40% of the incoming proton beam energy, proton vertex factorization seems to hold. JAM imposes a stricter requirement of $x_{L}>0.8$, corresponding to pions with less than 20% of the incoming proton beam energy. Applying this requirement, the accessible pion energies for the 10x275 GeV configuration are only 55 GeV and below. It should be remembered that only the region of $1\lesssim E_{\gamma}\lesssim 6~\mathrm{GeV}$ resides within the Low-$Q^{2}$ tagger acceptance, so the range of center-of-mass energies is only roughly $7\lesssim\sqrt{s_{\gamma\pi}}\lesssim 36\mathrm{~GeV}$. This range of energies is similar to existing and proposed fixed-target Drell–Yan experiments at Fermilab, CERN, and JPARC.
@@
Figure 13: Diagram of leading order Drell–Yan in $\gamma\pi^{+}$. This process can be accessed at the EIC by tagging a leading neutron and requiring a large rapidity gap between the neutron and the pion-remnant system $X_{\pi}$.
@@
The primary benefits of the $\gamma\pi$ Drell–Yan measurement at the EIC are to test key theoretical assumptions and to access previously unexplored regions of phase space in $Q^{2}$ and $x_{\pi}$. It is unlikely that the measurement precision at the EIC will compete with the precision of the leading-neutron DIS or pion Drell–Yan data. However, the universality of proton vertex factorization at collider energies can be tested for the first time in a process other than DIS, helping to validate the assumption that pion structure is accessible in leading-neutron processes more generally.
@@ VII Acknowledgments
I am grateful to the Sherpa team for implementing resolved photoproduction in their event generator and to P. Meinzinger in particular for help understanding and running Sherpa. Without their efforts this work would not have been possible.
@@ Appendix A Detector Resolutions and Acceptances
The strawman detector used for the reconstruction studies follows closely the design of ePIC. When handling charged particles, whichever detector resolution is best is the assumed resolution; i.e., for pions in the barrel, the tracking resolution is used instead of the hadronic calorimetry resolution. $K^{0}_{s}$ and $\Lambda$ are assumed to have the momentum resolution of the tracking. Neutrons and $K^{0}_{L}$ are produced in roughly equal numbers, so long-lived neutral hadrons are assigned the average of the two masses, i.e., 0.72 GeV. Particle identification of charged hadrons is assumed to be correct within the central detector acceptance.
@@
Table 4: Low-$Q^{2}$ tagger smearing and acceptance, applied to the scattered electron in the tagged sample.